Pabon Corporation makes one product. Budgeted unit sales for August and September are 11 00 and 12600 Units. respect el The ending finished goods inventory equals 40% of the follo ing months sales. The direct labor wage rate is $19.00 per hour. Each unit of finished goods requires 25 direct labor-hours. The estimated direct labor cost for August is closest to Multiplo Choice $389,000 555750 $29.250 $222.300

Pabon Corporation makes one product. Budgeted unit sales for August and September are 11 00 and 12600 Units. respect el The ending finished goods inventory equals 40% of the following month's sales. The direct labor wage rate is $19.00 per hour. Each unit of finished goods requires 25 direct labor-hours. The estimated direct labor cost for August is closest to 

Multiple Choice 

  • $389,000 

  • $55750 

  • $29.250 

  • $222.300
Answer #1

Calculate direct labour cost of august :

August production unit = 11100+(12600*40%)-(11100*40%) = 11700

Direct labour cost = 11700*2.5*19 = 555750

so answer is b) 555750
Answer #2

Calculate direct labor cost of august :

August production unit = 11100+(12600*40%)-(11100*40%) = 11700

Direct labour cost = 11700*2.5*19 = 555750

so answer is b) 555750



answered by: Pooja kasera